# Implicit Bias of SignGD and Adam on Multiclass Separable Data

### v1

@@ F.1 Exponential Loss
$\langle{\bm{A}},-\nabla\mathcal{L}_{\exp}({\bm{W}})\rangle=\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}\exp\left(-(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)\cdot\left(\bm{e}_{y_{i}}-\bm{e}_{c}\right)^{\top}{\bm{A}}\bm{h}_{i}\,.$
@@
from which we recognize that ${\mathcal{G}}_{\exp}({\bm{W}})=\mathcal{L}_{\exp}({\bm{W}})$ and the rest follows.
@@ F.2 PairLogLoss
The PairLogLoss loss (Wang et al., 2021b) is given as
@@
Lemma F.1 (Analogue of Lemma B.1 for PLL).
@@
$2B\cdot{\mathcal{G}}_{\rm{pll}}({\bm{W}})\geq{\left\|\nabla\mathcal{L}_{\rm{pll}}({\bm{W}})\right\|_{\rm{sum}}}\geq\gamma\cdot{\mathcal{G}}_{\rm{pll}}({\bm{W}})\,.$
@@
The lower bound follows immediately from (33) and expressing ${\left\|\nabla\mathcal{L}_{\rm{pll}}({\bm{W}})\right\|_{\rm{sum}}}=\max_{{\left\|{\bm{A}}\right\|_{\max}}\leq 1}\langle{\bm{A}},-\nabla\mathcal{L}_{\rm{pll}}({\bm{W}})\rangle$. The lower bound follows from triangle inequality applied to (32):
${\left\|\nabla\mathcal{L}_{\rm{pll}}({\bm{W}})\right\|_{\rm{sum}}}\leq\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}\left|f^{\prime}\left(-(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)\right|\|\bm{e}_{y_{i}}-\bm{e}_{c}\|_{1}\|\bm{h}_{i}\|_{1}\leq 2B\cdot{\mathcal{G}}({\bm{W}})\,.$
@@
$\bm{h}_{i}^{\top}\bm{\Delta}^{\top}(\bm{e}_{y_{i}}-\bm{e}_{c})(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}\bm{\Delta}\bm{h}_{i}$ $=\langle(\bm{e}_{y_{i}}-\bm{e}_{c})(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top},\bm{\Delta}\bm{h}_{i}\bm{h}_{i}^{\top}\bm{\Delta}^{T}\rangle\leq{\left\|(\bm{e}_{y_{i}}-\bm{e}_{c})(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}\right\|_{\rm{sum}}}{\left\|\bm{\Delta}\bm{h}_{i}\bm{h}_{i}^{\top}\bm{\Delta}^{T}\right\|_{\max}}$
$\leq\|\bm{e}_{y_{i}}-\bm{e}_{c})\|_{1}^{2}\cdot\|\bm{\Delta}\bm{h}_{i}\|_{\infty}^{2}$
$\leq 4\cdot\left({\left\|\bm{\Delta}\right\|_{\max}}\right)^{2}\cdot\|\bm{h}_{i}\|_{1}^{2}\leq 4B^{2}\left({\left\|\bm{\Delta}\right\|_{\max}}\right)^{2}\,.$
@@
$\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}f^{\prime\prime}\left((\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)\cdot\bm{h}_{i}^{\top}\bm{\Delta}^{\top}(\bm{e}_{y_{i}}-\bm{e}_{c})(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}\bm{\Delta}\bm{h}_{i}\leq 4B^{2}\cdot{\mathcal{G}}({\bm{W}})\cdot\left({\left\|\bm{\Delta}\right\|_{\max}}\right)^{2}\,.$
Finally, we verify PLL satisfies Lemma B.3.
Lemma F.2 (Analogue of Lemma B.3 for PLL).
@@
where the last line used $\log(1+x_{ic})\geq 0$. For (15), let $a=\log(1+x)>0$. The inequality becomes $e^{-a}\leq 1-a+a^{2}/2$, which holds for $a>0$ by the second-order Taylor expansion of $e^{-a}$ around 00.
@@
Lemma F.3 (Analogue of Lemma B.5 for PLL).
@@
$\frac{{\mathcal{G}}_{pll}({\bm{W}}+\psi\triangle{\bm{W}})}{{\mathcal{G}}_{pll}({\bm{W}})}\leq e^{2B\psi{\left\|\triangle{\bm{W}}\right\|_{\max}}}+2$
@@
Denote $x_{ic}^{{\bm{W}}}:=(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}\bm{h}_{i}$ and $x_{ic}^{{\bm{W}}^{\prime}}:=(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}({\bm{W}}+\psi\bm{\Delta}{\bm{W}})\bm{h}_{i}$, then we have for $i\in[n]$, $c\neq y_{i}$
$\frac{f^{\prime}(x_{ic}^{{\bm{W}}^{\prime}})}{f^{\prime}(x_{ic}^{{\bm{W}}})}=|\frac{f^{\prime}(x_{ic}^{{\bm{W}}^{\prime}})}{f^{\prime}(x_{ic}^{{\bm{W}}})}|\leq e^{|x_{ic}^{{\bm{W}}}-x_{ic}^{{\bm{W}}^{\prime}}|}+2$ $=e^{\psi|(\bm{e}_{c}-\bm{e}_{y_{i}})^{T}\bm{\Delta}{\bm{W}}\bm{h}_{i}|}+2=e^{\psi|\langle\bm{\Delta}{\bm{W}},(\bm{e}_{c}-\bm{e}_{y_{i}})\bm{h}_{i}^{T}\rangle|}+2$
$\leq e^{\psi\lVert\bm{\Delta}{\bm{W}}\rVert_{\max}{\left\|(\bm{e}_{c}-\bm{e}_{y_{i}})\bm{h}_{i}^{T}\right\|_{\rm{sum}}}}+2$
$=e^{\psi\lVert\bm{\Delta}{\bm{W}}\rVert_{\max}{\left\|\bm{e}_{c}-\bm{e}_{y_{i}}\right\|_{\rm{sum}}}{\left\|\bm{h}_{i}\right\|_{\rm{sum}}}}+2$
$\leq e^{2B\psi{\left\|\bm{\Delta}{\bm{W}}\right\|_{\max}}}+2.$
This leads to $\sum_{i\in[n]}\sum_{c\neq y_{i}}f^{\prime}(x_{ic}^{{\bm{W}}^{\prime}})\leq(e^{2B\psi{\left\|\bm{\Delta}{\bm{W}}\right\|_{\max}}}+2)\sum_{i\in[n]}\sum_{c\neq y_{i}}f^{\prime}(x_{ic}^{{\bm{W}}})$. Rearrange and using the definition of ${\mathcal{G}}_{pll}({\bm{W}})$, we obtain the desired. ∎
Lemma F.4 (Analogue of Lemma B.4 for PLL).
@@
Lemma F.5 (Analogue of Lemma B.2 for PLL).
@@
$|\mathcal{L}_{pll}({\bm{W}})-\mathcal{L}_{pll}({\bm{W}}_{0})|\leq 2B{\left\|{\bm{W}}-{\bm{W}}_{0}\right\|_{\max}}.$
@@
This lemma is a direct consequence of Lemma F.1 and be proved in the same way as Lemma B.2. ∎

### v4

@@ F.1 Exponential Loss
$\langle{\bm{A}},-\nabla\mathcal{L}_{\exp}({\bm{W}})\rangle=\frac{1}{n}\sum\nolimits_{i\in[n]}\sum\nolimits_{c\neq y_{i}}\exp\left(-(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)\cdot\left(\bm{e}_{y_{i}}-\bm{e}_{c}\right)^{\top}{\bm{A}}\bm{h}_{i}\,.$
@@
from which we recognize that ${\mathcal{G}}_{\exp}({\bm{W}})=\mathcal{L}_{\exp}({\bm{W}})$. Then, the proof follows similar steps as the CE loss.
@@ F.2 PairLogLoss
The PairLogLoss loss [wang2021rank4class] is given as
@@
Lemma 28 (Analogue of Lemma 16 for PLL).
@@
$2B\cdot{\mathcal{G}}_{\rm{pll}}({\bm{W}})\geq\lVert\nabla\mathcal{L}_{\rm{pll}}({\bm{W}})\rVert\geq\gamma\cdot{\mathcal{G}}_{\rm{pll}}({\bm{W}})\,.$
@@
The lower bound follows immediately from (29) and expressing $\lVert\nabla\mathcal{L}_{\rm{pll}}({\bm{W}})\rVert_{*}=\max_{\lVert{\bm{A}}\rVert\leq 1}\langle{\bm{A}},-\nabla\mathcal{L}_{\rm{pll}}({\bm{W}})\rangle$. The lower bound follows from triangle inequality applied to (28):
${\left\|\nabla\mathcal{L}_{\rm{pll}}({\bm{W}})\right\|_{\rm{sum}}}\leq\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}\left|f^{\prime}\left(-(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)\right|\|\bm{e}_{y_{i}}-\bm{e}_{c}\|_{1}\|\bm{h}_{i}\|_{1}\leq 2B\cdot{\mathcal{G}}({\bm{W}})\,$,
and use the relationships in (7), i.e. $\lVert\nabla\mathcal{L}_{\rm{pll}}({\bm{W}})\rVert\leq{\left\|\nabla\mathcal{L}_{\rm{pll}}({\bm{W}})\right\|_{\rm{sum}}}$ for any entry-wise or Schatten p-norm with $p\geq 1$. ∎
@@
$\bm{h}_{i}^{\top}\bm{\Delta}^{\top}(\bm{e}_{y_{i}}-\bm{e}_{c})(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}\bm{\Delta}\bm{h}_{i}$ $=\langle(\bm{e}_{y_{i}}-\bm{e}_{c})(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top},\bm{\Delta}\bm{h}_{i}\bm{h}_{i}^{\top}\bm{\Delta}^{T}\rangle$
$\leq{\left\|(\bm{e}_{y_{i}}-\bm{e}_{c})(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}\right\|_{\rm{sum}}}{\left\|\bm{\Delta}\bm{h}_{i}\bm{h}_{i}^{\top}\bm{\Delta}^{T}\right\|_{\max}}$
$\leq\|\bm{e}_{y_{i}}-\bm{e}_{c})\|_{1}^{2}\cdot(\|\bm{\Delta}\bm{h}_{i}\|_{\infty})^{2}$
$\leq 4\cdot({\left\|\bm{\Delta}\right\|_{\max}})^{2}\cdot\|\bm{h}_{i}\|_{1}^{2}\leq 4B^{2}({\left\|\bm{\Delta}\right\|_{\max}})^{2}\,$
$\leq 4B^{2}\lVert\bm{\Delta}\rVert^{2}.$
@@
$\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}f^{\prime\prime}\left((\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)\cdot\bm{h}_{i}^{\top}\bm{\Delta}^{\top}(\bm{e}_{y_{i}}-\bm{e}_{c})(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}\bm{\Delta}\bm{h}_{i}\leq 4B^{2}\cdot{\mathcal{G}}({\bm{W}})\cdot\left(\lVert\bm{\Delta}\rVert\right)^{2}\,.$
Finally, we verify PLL satisfies Lemma 18.
Lemma 29 (Analogue of Lemma 18 for PLL).
@@
where the last line used $\log(1+x_{ic})\geq 0$. For (19), let $a=\log(1+x)>0$. The inequality becomes $e^{-a}\leq 1-a+a^{2}/2$, which holds for $a>0$ by the second-order Taylor expansion of $e^{-a}$ around 0.
@@
Lemma 30 (Analogue of Lemma 20 for PLL).
@@
$\frac{{\mathcal{G}}_{pll}({\bm{W}}-\psi\eta\bm{\Delta})}{{\mathcal{G}}_{pll}({\bm{W}})}\leq e^{2B\psi\lVert\triangle{\bm{W}}\rVert}+2$
@@
Denote $x_{ic}^{{\bm{W}}}:=(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}\bm{h}_{i}$ and $x_{ic}^{{\bm{W}}^{\prime}}:=(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}({\bm{W}}-\psi\eta\bm{\Delta})\bm{h}_{i}$, then we have for $i\in[n]$, $c\neq y_{i}$
$\frac{f^{\prime}(x_{ic}^{{\bm{W}}^{\prime}})}{f^{\prime}(x_{ic}^{{\bm{W}}})}=|\frac{f^{\prime}(x_{ic}^{{\bm{W}}^{\prime}})}{f^{\prime}(x_{ic}^{{\bm{W}}})}|\leq e^{|x_{ic}^{{\bm{W}}}-x_{ic}^{{\bm{W}}^{\prime}}|}+2$ $=e^{\psi\eta|(\bm{e}_{c}-\bm{e}_{y_{i}})^{T}\bm{\Delta}\bm{h}_{i}|}+2=e^{\psi\eta|\langle\bm{\Delta},(\bm{e}_{c}-\bm{e}_{y_{i}})\bm{h}_{i}^{T}\rangle|}+2$
$\leq e^{\psi\eta\lVert\bm{\Delta}\rVert_{\max}{\left\|(\bm{e}_{c}-\bm{e}_{y_{i}})\bm{h}_{i}^{T}\right\|_{\rm{sum}}}}+2$
$=e^{\psi\eta\lVert\bm{\Delta}\rVert_{\max}{\left\|\bm{e}_{c}-\bm{e}_{y_{i}}\right\|_{\rm{sum}}}{\left\|\bm{h}_{i}\right\|_{\rm{sum}}}}+2$
$\leq e^{2B\psi\eta{\left\|\bm{\Delta}\right\|_{\max}}}+2.$
This leads to $\sum_{i\in[n]}\sum_{c\neq y_{i}}f^{\prime}(x_{ic}^{{\bm{W}}^{\prime}})\leq(e^{2B\psi{\left\|\bm{\Delta}{\bm{W}}\right\|_{\max}}}+2)\sum_{i\in[n]}\sum_{c\neq y_{i}}f^{\prime}(x_{ic}^{{\bm{W}}})$. Rearrange and using the definition of ${\mathcal{G}}_{pll}({\bm{W}})$ and relationships in (7), we obtain the desired. ∎
Lemma 31 (Analogue of Lemma 19 for PLL).
@@
Lemma 32 (Analogue of Lemma 17 for PLL).
@@
$|\mathcal{L}_{pll}({\bm{W}})-\mathcal{L}_{pll}({\bm{W}}_{0})|\leq 2B\lVert{\bm{W}}-{\bm{W}}_{0}\rVert.$
@@
This lemma is a direct consequence of Lemma 28 and can be proved in the same way as Lemma 17. ∎